OpenStudy (anonymous):

Which of the following equations does not require two steps to solve? negative eight q minus six equals ten negative nine plus s equals negative one four h minus twenty-two equals negative seven seventeen minus f equals 4

OpenStudy (anonymous):

SOMEONE PLZ HELP

OpenStudy (anonymous):

the last one is 17 - f = 4 solve it and count how many steps it takes

OpenStudy (anonymous):

Which of the following equations does not require two steps to solve? negative seven b minus thirteen equals one sixteen minus r equals eight negative thirty-two plus three k equals negative twenty-six w minus twenty-two equals negative fifteen

OpenStudy (anonymous):

write them all out as equations. solve them, count the steps it took you.

OpenStudy (anonymous):

Thats what the question is asking you to do. Which bit cant you do?

OpenStudy (anonymous):

Which property would you use to solve the equation w divided by seven equals zero point one four for w? Addition Property of Equality Multiplication Property of Equality Subtraction Property of Equality Division Property of Equality

OpenStudy (anonymous):

w/7 = 4 solve it. How did you solve it, thats the property.

OpenStudy (anonymous):

I AM ONLY IN THE SEVENTH THAT IS Y I AM ASK AND I DON'T WANT TO FAIL Y CANT YOU HELP

OpenStudy (anonymous):

okay so w/7 = 4 we want to know what w equals so we want w = (something) read as "w = something" w/7 = 4, how can we get the w on its own we need to get rid of the 7 it is being divided by. what is the opposite of division, what is it that "undoes" dividing?